ganeshie8 (ganeshie8):

Complete the following little proof showing that y^3 = x^2 + 2 has no solutions over integers other than x=5, y=3

ganeshie8 (ganeshie8):

\(\begin{align}y^3 &= x^2+2\\~\\ &=x^2 - (i\sqrt{2})^2\\~\\ &=(x+i\sqrt{2})(x-i\sqrt{2}) \end{align}\) Each of the right hand side factors must be a perfect cube since the left hand side is a perfect cube. In particular, we take : \[\begin{align}x+i\sqrt{2} &= (a+ib\sqrt{2})^3\\~\\ &=\cdots \end{align} \]

OpenStudy (baru):

x=3 y=2 is not a solution.. ?

ILovePuppiesLol (ilovepuppieslol):

dang :( doing stuff i don't know ;-;

ganeshie8 (ganeshie8):

sorry il update, it is x=5, y=3

ganeshie8 (ganeshie8):

you do know these stuff @ILovePuppiesLol i have just phrased the question in a way that you wont understand easily.. ;)

ILovePuppiesLol (ilovepuppieslol):

can you clear it up for me so i don't sit here like a pregnant koala trying to swim, please :)

ganeshie8 (ganeshie8):

In simple words, here is the problem : ``` Find all the whole numbers that satisfy the equation y^3 = x^2 + 2 ```

ILovePuppiesLol (ilovepuppieslol):

ohhh

ganeshie8 (ganeshie8):

plugin x = 5, y = 3 and convince yourself that it is indeed a solution

ILovePuppiesLol (ilovepuppieslol):

yes, it is, wait what are we trying to figure out again, as many solutions as possible?

ganeshie8 (ganeshie8):

I challenge you to find just one more solution

ILovePuppiesLol (ilovepuppieslol):

will substitution work, or will it take me about a year

ILovePuppiesLol (ilovepuppieslol):

x=-5 xD

ILovePuppiesLol (ilovepuppieslol):

y=3

ganeshie8 (ganeshie8):

Very very clever puppy!

ILovePuppiesLol (ilovepuppieslol):

:O I did it

ganeshie8 (ganeshie8):

Here is your medal for being alert

ILovePuppiesLol (ilovepuppieslol):

yay!

ILovePuppiesLol (ilovepuppieslol):

what do we do now?

ganeshie8 (ganeshie8):

Try finding another solution

ILovePuppiesLol (ilovepuppieslol):

another one? UGHHH

ILovePuppiesLol (ilovepuppieslol):

im stuck

ILovePuppiesLol (ilovepuppieslol):

y^3-x^2 = 2 is basically what your equation says right?

ganeshie8 (ganeshie8):

Yes, im deleting the graph it is making the thread slow..

OpenStudy (bobo-i-bo):

I did it! Followed through and got \(x=\pm 5, y=3\) as the solution. The way they go complex in this proof is pretty new to me, it's interesting! :)

ganeshie8 (ganeshie8):

Awesome! It must have felt amazing finishing off the proof :D

OpenStudy (bobo-i-bo):

Haha, I'm afraid the cool bit was done by you, it was great to learn about the method but the proof was relatively trivial after that ;P

OpenStudy (baru):

solution please :| is there any other way to do it?

OpenStudy (baru):

without complex numbers...

ILovePuppiesLol (ilovepuppieslol):

solve \[y^3-x^2 = 2\] @baru

OpenStudy (mww):

I find you can just graph it, that easily shows you there can be at most two real solutions. lol

OpenStudy (anonymous):

For every value of x there is a corresponding value of y So there are infinitely many real solutions